OpenStudy (anonymous):

solve the system by graphing. -3x-y=-10 4x-4y=8 open study keeps freezing on me and will not load right every time, if you have a facebook, kik, please post what you have and username below.

OpenStudy (triciaal):

one approach For each equation find when x = 0 what is y and when y = 0 what is x you will have the minimum 2 points to draw the line Draw the 2 lines and the solution is where they meet.

OpenStudy (anonymous):

I don't know how to put this im in a slow math and im in cyber school they don't explain how to do anything so I have no clue how to find x or any of this that's why I posted it on here

OpenStudy (anonymous):

and do you have a facebook because any time I write a reply it freezes and the box is invisible, and then my computer freezes,

OpenStudy (triciaal):

for the 2nd eqn 4x-4y=8 you can divide by 4 first when x = 0 -4y = 8 so y = -2 point (0, -2) when y = 0 4x = 8 so x = 2 point (2, 0)